# Crossed products by compact group actions with the weak tracial Rokhlin property

### v1

@@ Abstract.
In this paper, we introduce compact group actions with the weak tracial Rokhlin property. This concept simultaneously generalizes finite group actions with the weak tracial Rokhlin property and compact group actions with the tracial Rokhlin property (in the sense of the Elliott program). Under this framework, we prove that simplicity, pure infiniteness, tracial $\mathcal{Z}$-stability and the combination of nuclearity and $\mathcal{Z}$-stability can be transferred from the original algebra to the crossed product. We also show that the radius of comparison of the fixed point algebra does not exceed that of the original algebra. Furthermore, we discuss the relationship between our definitions and natural generalizations of the finite group case in non-Elliott program settings. Finally, we provide a nontrivial example of a compact group action with the weak tracial Rokhlin property with comparison: an action of $(S_{2})^{\mathbb{N}}$ on the Jiang-Su algebra $\mathcal{Z}$. Since $\mathcal{Z}$ contains no nontrivial projections, this action does not possess the tracial Rokhlin property.
@@ 1. Introduction
In this paper, We introduce compact group actions with the weak tracial Rokhlin property. This concept simultaneously generalizes finite group actions with the weak tracial Rokhlin property and compact group actions with the tracial Rokhlin property (in the sense of the Elliott program). As in the case of finite groups, it is impossible for a compact group action on a C*-algebra containing no non-trivial projections (such as the Jiang-Su algebra $\mathcal{Z}$) to have the tracial Rokhlin property. Therefore, this generalization has practical research value. Under this concept, we prove that simplicity, pure infiniteness, tracial $\mathcal{Z}$-stability, and the combination of amenability and $\mathcal{Z}$-stability can be transferred from the original algebra to the crossed product algebra. We also prove that the radius of comparison of the fixed point algebra does not exceed that of the original algebra. It is worth mentioning that during the research process, we did not directly study the crossed product algebra, but instead employed the so-called “averaging process” introduced by Gardella[16]. That is, we first study how structural properties transfer from the original algebra to the fixed point algebra, and then utilize the fact that the fixed point algebra and the crossed product algebra are Morita equivalent (and hence stably isomorphic in the separable case) to obtain the structural properties of the crossed product algebra. Simultaneously, we discuss the relationship between our definition and the natural generalization of the finite group case in a non-Elliott program sense. Finally, we provide a non-trivial example of a compact group action with the weak tracial Rokhlin property and finite radius of comparison. This is an action of $(S_{2})^{\mathbb{N}}$ on the Jiang-Su algebra $\mathcal{Z}$. Since the Jiang-Su algebra $\mathcal{Z}$ contains no non-trivial projections, this action does not have the tracial Rokhlin property.
@@ 2.2. Cuntz comparison
Below we give some well-known facts about Cuntz comparison. (1) is contained in [30, Proposition 2.6] (and also appears in a slightly different form in the earlier [43, Proposition 2.4]). (2) is [41, Lemma 1.7]. (3) is straightforward (and is also [30, Lemma 2.5(i)]). (4) is [41, Corollary 1.6]. (5) is [30, Lemma 2.8(ii)]. (6) is [30, Lemma 2.8(iii)]. (7) is [30, Lemma 2.9]. (8) is [31, Lemma 2.4(2)]. For (9), it is straightforward to verify that $a$ and $f(a)$ generate the same hereditary subalgebra in $A$. (10) is contained in the discussion following [30, Definition 2.3].
@@
[41, Lemma 2.4] Let $A$ be a simple C*-algebra which is not of type I. Let $a\in A$ be a nonzero positive element, and $l$ a positive integer. Then there exist $l$ mutually orthogonal nonzero positive elements $b_{1},\ldots,b_{l}\in A$ such that $b_{1}\sim_{A}\cdots\sim_{A}b_{l}$, and $b_{1}+\cdots+b_{l}\in\overline{aAa}$.
@@
[41, Lemma 1.5] Let $A$ be a C*-algebra, let $a,b\in A_{+}$, and let $\alpha,\beta\geq 0$. Then
@@
The following lemma is known as Kirchberg’s Slice Lemma, see [42, Lemma 4.1.9].
@@
Below we introduce the strict comparison of positive elements and the radius of comparison.
@@
    For $a\in(A\otimes K)_{+}$, $\langle a\rangle\in\mathrm{Cu}(A)$ and $\langle a\rangle\in\mathrm{W}(A)$, we use the same notation for the corresponding functional. From part of the proof of [10, Proposition 4.2], it follows that $d_{\tau}$ is well-defined on $\mathrm{Cu}(A)$ and $\mathrm{W}(A)$.
@@
    According to [7, Theorem I.2.2], for any $\tau\in\mathrm{QT}(A)$, $d_{\tau}$ is a lower semicontinuous function on $A$. Consequently, for any $a\in A_{+}\setminus\{0\}$ we have
@@
By replacing normalised 2-quasitraces with densely defined 2-quasitraces, the strict comparison property for positive elements can also be defined for non-unital C*-algebras (see [34]). Recall that an ordered semigroup $S$ is called almost unperforated if for any $x,y\in S$ and $n\in\mathbb{N}$ satisfying $(n+1)x\leq ny$, we have $x\leq y$. By [47, Lemma 6.1], for a simple C*-algebra, having strict comparison of positive elements is equivalent to its Cuntz semigroup being almost unperforated.
@@
[48, Definition 6.1] Let $A$ be a unital stably finite C*-algebra. Let $r\in[0,\infty)$. We say that $A$ has $r$-comparison if whenever positive elements $a,b$ in $M_{\infty}(A)$ satisfy $d_{\tau}(a)+r<d_{\tau}(b)$ for all $\tau\in\mathrm{QT}(A)$, then $a\precsim b$. Define the radius of comparison of $A$ as
@@
The above definition is equivalent to defining strict comparison and the radius of comparison using $\mathrm{W}(A)$; using $\mathrm{Cu}(A)$ yields the same result (see [41, Proposition 6.12]). If $A$ is simple and has finite radius of comparison, then the above infimum is attained, meaning $A$ has $\mathrm{rc}(A)$-comparison (see [48, Proposition 6.3] and [41, Lemma 1.23]).
@@ 2.3. Completely positive order zero map
[53, Definition 2.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. We say $\varphi$ has order zero if for every $a,b\in A$, we have $\varphi(a)\perp\varphi(b)$ whenever $a\perp b$.
@@
[22, Theorem 2.2, Corollary 4.8] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. Then $\varphi$ has order zero if and only if for any $a,b,c\in A$, we have
@@
In the case where $A$ has a unit, the above theorem follows easily from the correspondence between completely positive order zero maps and homomorphisms given below; for the general case, see the relevant proof in [22].
@@
[53, Theorem 3.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map. Let $C\colon C^{*}(\varphi(A))\subset B$.
@@
[53, Corollary 4.1] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive contractive order zero map. Then, the map defined by $\rho_{\varphi}(\mathrm{id}_{(0,1]}\otimes a)\colon=\varphi(a)$ induces a *-homomorphism $\rho_{\varphi}\colon C_{0}((0,1])\otimes A\to B$.
@@
[53, Corollary 4.2] Let $\varphi\colon A\to B$ be a completely positive contractive order zero map, and let $f\in C_{0}((0,1])$ be a positive function. Let $C$, $h$, and $\pi_{\varphi}$ be as in Theorem 2.16. Then, the map
@@
[53, Corollary 4.3] Let $A,B,C$ and $D$ be C*-algebras and let $\varphi\colon A\to B$ and $\psi\colon C\to D$ be completely positive contractive order zero maps.
@@
[53, Corollary 4.4] Let $A$ and $B$ be C*-algebras, $\varphi\colon A\to B$ a completely positive order zero map, and $\tau$ a trace on $B$.
@@
[53, Corollary 4.5] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map.
@@ 2.4. Maps under group actions
[28, Definition 1.3] Let $G$ be a compact group, and let $A$, $B$ be C*-algebras. Let $\alpha\colon G\to\mathrm{Aut}(A)$ and $\gamma\colon G\to\mathrm{Aut}(B)$ be actions of $G$ on $A$ and $B$, respectively. Let $F\subseteq A$, $S\subseteq B$, and $\varepsilon>0$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(F,S,\varepsilon)$-approximately centrally equivariant multiplicative map if:
@@
By ignoring the group action and omitting condition 3, we obtain the definition of a general $(F,S,\varepsilon)$-approximately centrally multiplicative map.
@@
[36, Definition 1.4] Let $A$ and $B$ be C*-algebras, and let $F\subseteq A$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(n,F,\varepsilon)$-approximately multiplicative map if for any $m\in\{1,2,\ldots,n\}$ and any $x_{1},x_{2},\ldots,x_{m}\in F$, we always have
@@
If $S\subseteq B$ is also given, then $\varphi$ is said to be an $(n,F,S,\varepsilon)$-approximately centrally multiplicative map if it also satisfies $\|\varphi(x)a-a\varphi(x)\|<\varepsilon$ for all $x\in F$ and all $a\in S$.
@@
Replacing approximate multiplicativity with strong approximate order zero yields the following definition.
@@
Let $G$ be a compact group, let $A$ be a unital C*-algebra, and $B$ a C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ and $\gamma\colon G\to\mathrm{Aut}(B)$ be actions of $G$ on $A$ and $B$, respectively. Let $F\subseteq A$, $S\subseteq B$, and $\varepsilon>0$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(F,S,\varepsilon)$-approximately centrally equivariant strongly order zero map if:
@@ 3. The weak tracial Rokhlin property with comparison
[6, Definition 3.2] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that $\alpha$ has the weak tracial Rokhlin property if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exist mutually orthogonal contractive positive elements $(d_{g})_{g\in G}\in A$ with $d=\sum_{g\in G}d_{g}$ such that
@@
[5, Lemma 2.14] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property if and only if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exists a positive element $d\in A^{\alpha}$ with $\|d\|=1$ and contractive positive elements $(d_{g})_{g\in G}\in A$ such that
@@
[36, Definition 2.4] Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the tracial Rokhlin property with comparison if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a projection $p\in A^{\alpha}$ and a unit-preserving completely positive map $\psi\colon C(G)\to pAp$ such that
@@
The approximate multiplicativity requirement in Definition 3.3 implies that $A$ must contain sufficiently many projections. However, for C*-algebras without non-trivial projections (such as the Jiang-Su algebra $\mathcal{Z}$), this means no such group action can exist on them. Approximate order zero maps impose no such restriction and can also approximately preserve orthogonality. Synthesizing the above two definitions, we now give the definition of the weak tracial Rokhlin property with comparison for compact group actions.
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central equivariant approximate order zero map.
@@
Definition 3.4 remains valid under the following modification: if we assume $F$ and $S$ are compact sets instead of finite sets, and require $\psi\colon C(G)\to\overline{dAd}$ to be fully equivariant.
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central approximate order zero map.
@@
The proof is essentially the same as [36, Lemma 2.9], so we omit it. ∎
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central approximate order zero map.
@@
The proof is essentially the same as [12, Lemma 3.9], so we omit it. ∎
When $A$ is finite, condition (6) of Definition 3.4, condition (6) of Lemma 3.6, and condition (6) of Lemma 3.7 are redundant. To prove this, we need a technique from [41, Lemma 4.5].
@@
[6, Lemma 2.6] Let $A$ be a unital C*-algebra, let $a,g\in A_{+}$ with $0\leq a,g\leq 1$, and let $\varepsilon_{1},\varepsilon_{2}\geq 0$. Then
@@
The proof is essentially the same as [41, Lemma 2.9]; we simply replace [41, Lemma 1.8] with Lemma 3.8. ∎
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central equivariant approximate order zero map.
@@
The proof is essentially the same as [41, Lemma 4.5]; we simply replace [41, Lemma 2.9] with Lemma 3.9. ∎
When $A$ is separable, we characterize the weak tracial Rokhlin property with comparison using the language of central sequences. We begin with a definition.
@@
By (7), (8), and (10), we can choose $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, we have
$\|\gamma_{n}(1)\gamma_{n}(f_{1}f_{2})-\gamma_{n}(f_{1})\gamma_{n}(f_{2})\|<\varepsilon\quad\mbox{for all }f_{1},f_{2}\in S$
and
$\|\gamma_{n}(f)a-a\gamma_{n}(f)\|<\varepsilon\quad\mbox{for all }a\in F\mbox{ and }f\in S.$
Since $1-d\precsim_{\alpha-p.w.-A}x$, we can choose $n_{1}\in\mathbb{N}$ such that for all $n\geq n_{1}$, we have
$(1-h_{n}-\varepsilon)_{+}\precsim_{A}x.$
Similarly, since $1-d\precsim_{p.w.-A^{\alpha}}y$, we can choose $n_{2}\in\mathbb{N}$ such that for all $n\geq n_{2}$, we have
$(1-h_{n}-\varepsilon)_{+}\precsim_{A^{\alpha}}y.$
Similarly, since $1-d\precsim_{p.w.-A^{\alpha}}d$, we can choose $n_{3}\in\mathbb{N}$ such that for all $n\geq n_{3}$, we have
$(1-h_{n}-\varepsilon)_{+}\precsim_{A^{\alpha}}h_{n}.$
Also, we can choose $n_{4}\in\mathbb{N}$ such that for all $n\geq n_{4}$, we have
$\|h_{n}xh_{n}\|>1-\varepsilon.$
Set $N=\max(n_{0},n_{1},n_{2},n_{3},n_{4})$, choose $\psi=\gamma_{N}$ and $d=h_{N}$, then all conditions are satisfied.
@@
    $\psi_{n}$ is an ($S_{n}$$F_{n}$$\frac{1}{n}$)-approximately central approximate order zero map.
@@
We now introduce an “averaging process” introduced by Gardella [16], which allows us to transfer certain properties from the original algebra to the fixed point algebra.
@@
[16, Lemma 4.2] Let $A$ be a unital C*-algebra, let $G$ be a compact group, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be a continuous action. Let $\varphi\colon C(G,A)\to C(G,A)$ be the averaging process defined above, and let $\gamma\colon G\to\mathrm{Aut}(C(G,A))$ be the diagonal action. Given $\varepsilon>0$ and a compact set $F\subseteq C(G,A)$, there exist $\delta>0$, a finite set $K\subseteq G$, and continuous functions $f_{k}\in C(G)$ for $k\in K$ such that
@@
The following proof references [16, Proposition 4.3].

### v2

@@ Abstract.
In this paper, we introduce compact group actions with the weak tracial Rokhlin property. This concept generalizes both the finite group actions with the weak tracial Rokhlin property and the compact group actions with the tracial Rokhlin property on classifiable C*-algebras (in the sense of the Elliott program). Under this framework, we prove that simplicity, pure infiniteness, tracial $\mathcal{Z}$-stability and the combination of nuclearity and $\mathcal{Z}$-stability can be transferred from the original algebra to the crossed product. We also show that the radius of comparison of the fixed point algebra does not exceed that of the original algebra. Furthermore, we discuss the relationship between our definitions and natural generalizations of the finite group case in non-Elliott program settings. Finally, we provide a nontrivial example of a compact group action with the weak tracial Rokhlin property with comparison: an action of $(S_{2})^{\mathbb{N}}$ on the Jiang-Su algebra $\mathcal{Z}$. Since $\mathcal{Z}$ contains no nontrivial projections, this action does not possess the tracial Rokhlin property.
@@ 1. Introduction
In this paper, We introduce compact group actions with the weak tracial Rokhlin property. This concept generalizes both the finite group actions with the weak tracial Rokhlin property and the compact group actions with the tracial Rokhlin property on classifiable C*-algebras (in the sense of the Elliott program). As in the case of finite groups, it is impossible for a compact group action on a C*-algebra containing no non-trivial projections (such as the Jiang-Su algebra $\mathcal{Z}$) to have the tracial Rokhlin property. Therefore, this generalization has practical research value. Under this concept, we prove that simplicity, pure infiniteness, tracial $\mathcal{Z}$-stability, and the combination of amenability and $\mathcal{Z}$-stability can be transferred from the original algebra to the crossed product algebra. We also prove that the radius of comparison of the fixed point algebra does not exceed that of the original algebra. It is worth mentioning that during the research process, we did not directly study the crossed product algebra, but instead employed the so-called “averaging process” introduced by Gardella in [17]. That is, we first study how structural properties transfer from the original algebra to the fixed point algebra, and then utilize the fact that the fixed point algebra and the crossed product algebra are Morita equivalent (and hence stably isomorphic in the separable case) to obtain the structural properties of the crossed product algebra. Simultaneously, we discuss the relationship between our definition and the natural generalization of the finite group case in a non-Elliott program sense. Finally, we provide a non-trivial example of a compact group action with the weak tracial Rokhlin property and finite radius of comparison. This is an action of $(S_{2})^{\mathbb{N}}$ on the Jiang-Su algebra $\mathcal{Z}$. Since the Jiang-Su algebra $\mathcal{Z}$ contains no non-trivial projections, this action does not have the tracial Rokhlin property.
@@ 2.2. Cuntz comparison
Below we give some well-known facts about Cuntz comparison. (1) is contained in [31, Proposition 2.6] (and also appears in a slightly different form in the earlier [45, Proposition 2.4]). (2) is [43, Lemma 1.7]. (3) is straightforward (and is also [31, Lemma 2.5(i)]). (4) is [43, Corollary 1.6]. (5) is [31, Lemma 2.8(ii)]. (6) is [31, Lemma 2.8(iii)]. (7) is [31, Lemma 2.9]. (8) is [32, Lemma 2.4(2)]. For (9), it is straightforward to verify that $a$ and $f(a)$ generate the same hereditary subalgebra in $A$. (10) is contained in the discussion following [31, Definition 2.3].
@@
[43, Lemma 2.4] Let $A$ be a simple C*-algebra which is not of type I. Let $a\in A$ be a nonzero positive element, and $l$ a positive integer. Then there exist $l$ mutually orthogonal nonzero positive elements $b_{1},\ldots,b_{l}\in A$ such that $b_{1}\sim_{A}\cdots\sim_{A}b_{l}$, and $b_{1}+\cdots+b_{l}\in\overline{aAa}$.
@@
[43, Lemma 1.5] Let $A$ be a C*-algebra, let $a,b\in A_{+}$, and let $\alpha,\beta\geq 0$. Then
@@
The following lemma is known as Kirchberg’s Slice Lemma, see [44, Lemma 4.1.9].
@@
Below we introduce the strict comparison of positive elements and the radius of comparison. The following definition of tracial state is from [34], and the definition of 2-quasitrace is from [8].
@@
    For $a\in(A\otimes K)_{+}$, $\langle a\rangle\in\mathrm{Cu}(A)$ and $\langle a\rangle\in\mathrm{W}(A)$, we use the same notation for the corresponding functional. From part of the proof of [11, Proposition 4.2], it follows that $d_{\tau}$ is well-defined on $\mathrm{Cu}(A)$ and $\mathrm{W}(A)$.
@@
    According to [8, Theorem I.2.2], for any $\tau\in\mathrm{QT}(A)$, $d_{\tau}$ is a lower semicontinuous function on $A$. Consequently, for any $a\in A_{+}\setminus\{0\}$ we have
@@
By replacing normalised 2-quasitraces with densely defined 2-quasitraces, the strict comparison property for positive elements can also be defined for non-unital C*-algebras (see [35]). Recall that an ordered semigroup $S$ is called almost unperforated if for any $x,y\in S$ and $n\in\mathbb{N}$ satisfying $(n+1)x\leq ny$, we have $x\leq y$. By [49, Lemma 6.1], for a simple C*-algebra, having strict comparison of positive elements is equivalent to its Cuntz semigroup being almost unperforated.
@@
[50, Definition 6.1] Let $A$ be a unital stably finite C*-algebra. Let $r\in[0,\infty)$. We say that $A$ has $r$-comparison if whenever positive elements $a,b$ in $M_{\infty}(A)$ satisfy $d_{\tau}(a)+r<d_{\tau}(b)$ for all $\tau\in\mathrm{QT}(A)$, then $a\precsim b$. Define the radius of comparison of $A$ as
@@
The above definition is equivalent to defining strict comparison and the radius of comparison using $\mathrm{W}(A)$; using $\mathrm{Cu}(A)$ yields the same result (see [43, Proposition 6.12]). If $A$ is simple and has finite radius of comparison, then the above infimum is attained, meaning $A$ has $\mathrm{rc}(A)$-comparison (see [50, Proposition 6.3] and [43, Lemma 1.23]).
@@ 2.3. Completely positive order zero map
[55, Definition 2.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. We say $\varphi$ has order zero if for every $a,b\in A$, we have $\varphi(a)\perp\varphi(b)$ whenever $a\perp b$.
@@
[23, Theorem 2.2, Corollary 4.8] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. Then $\varphi$ has order zero if and only if for any $a,b,c\in A$, we have
@@
In the case where $A$ has a unit, the above theorem can be easily derived from the decomposition theorem for completely positive order zero maps; for the general case, one may refer to the relevant proof in [23].
@@
[55, Theorem 3.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map. Let $C\colon C^{*}(\varphi(A))\subset B$.
@@
[55, Corollary 4.1] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive contractive order zero map. Then, the map defined by $\rho_{\varphi}(\mathrm{id}_{(0,1]}\otimes a)\colon=\varphi(a)$ induces a *-homomorphism $\rho_{\varphi}\colon C_{0}((0,1])\otimes A\to B$.
@@
[55, Corollary 4.2] Let $\varphi\colon A\to B$ be a completely positive contractive order zero map, and let $f\in C_{0}((0,1])$ be a positive function. Let $C$, $h$, and $\pi_{\varphi}$ be as in Theorem 2.16. Then, the map
@@
[55, Corollary 4.3] Let $A,B,C$ and $D$ be C*-algebras and let $\varphi\colon A\to B$ and $\psi\colon C\to D$ be completely positive contractive order zero maps.
@@
[55, Corollary 4.4] Let $A$ and $B$ be C*-algebras, $\varphi\colon A\to B$ a completely positive order zero map, and $\tau$ a trace on $B$.
@@
[55, Corollary 4.5] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map.
@@ 2.4. Maps under group actions
[29, Definition 1.3] Let $G$ be a compact group, and let $A$, $B$ be C*-algebras. Let $\alpha\colon G\to\mathrm{Aut}(A)$ and $\gamma\colon G\to\mathrm{Aut}(B)$ be actions of $G$ on $A$ and $B$, respectively. Let $F\subseteq A$, $S\subseteq B$, and $\varepsilon>0$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(F,S,\varepsilon)$-approximately equivariant central multiplicative map if:
@@
By ignoring the group action and omitting condition (3), we obtain the definition of a general $(F,S,\varepsilon)$-approximately central multiplicative map.
@@
[37, Definition 1.4] Let $A$ and $B$ be C*-algebras, and let $F\subseteq A$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(n,F,\varepsilon)$-approximately multiplicative map if for any $m\in\{1,2,\ldots,n\}$ and any $x_{1},x_{2},\ldots,x_{m}\in F$, we always have
@@
If $S\subseteq B$ is also given, then $\varphi$ is said to be an $(n,F,S,\varepsilon)$-approximately central multiplicative map if it also satisfies $\|\varphi(x)a-a\varphi(x)\|<\varepsilon$ for all $x\in F$ and all $a\in S$.
@@
Replacing approximate multiplicativity with strongly approximate order zero yields the following definition.
@@
Let $G$ be a compact group, let $A$ be a unital C*-algebra, and $B$ a C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ and $\gamma\colon G\to\mathrm{Aut}(B)$ be actions of $G$ on $A$ and $B$, respectively. Let $F\subseteq A$, $S\subseteq B$, and $\varepsilon>0$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(F,S,\varepsilon)$-approximately central equivariant strongly order zero map if:
@@ 3. The weak tracial Rokhlin property with comparison
[7, Definition 3.2] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that $\alpha$ has the weak tracial Rokhlin property if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exist mutually orthogonal contractive positive elements $(d_{g})_{g\in G}\in A$ with $d=\sum_{g\in G}d_{g}$ such that
@@
[6, Lemma 2.14] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property if and only if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exists a positive element $d\in A^{\alpha}$ with $\|d\|=1$ and contractive positive elements $(d_{g})_{g\in G}\in A$ such that
@@
[37, Definition 2.4] Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the tracial Rokhlin property with comparison if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a projection $p\in A^{\alpha}$ and a unital completely positive map $\psi\colon C(G)\to pAp$ such that
@@
The approximate multiplicativity requirement in Definition 3.3 implies that $A$ must contain sufficiently many projections. However, for C*-algebras without non-trivial projections (such as the Jiang-Su algebra $\mathcal{Z}$), this means no such group action can exist on them. Strongly order zero maps impose no such restriction and can also approximately preserve orthogonality. Synthesizing the above two definitions, we now give the definition of the weak tracial Rokhlin property with comparison for compact group actions.
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central equivariant strongly order zero map.
@@
Definition 3.4 remains valid under the following modification: if we assume $F$ and $S$ are compact sets instead of finite sets, and require $\psi\colon C(G)\to\overline{dAd}$ to be exactly equivariant.
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central strongly order zero map.
@@
The proof is essentially the same as [37, Lemma 2.9], so we omit it. ∎
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central strongly order zero map.
@@
The proof is essentially the same as [13, Lemma 3.9], so we omit it. ∎
When $A$ is finite, condition (6) of Definition 3.4, condition (6) of Lemma 3.6, and condition (6) of Lemma 3.7 are redundant. To prove this, we need a technique from [43, Lemma 4.5].
@@
[7, Lemma 2.6] Let $A$ be a unital C*-algebra, let $a,g\in A_{+}$ with $0\leq a,g\leq 1$, and let $\varepsilon_{1},\varepsilon_{2}\geq 0$. Then
@@
The proof is essentially the same as [43, Lemma 2.9]; we simply replace [43, Lemma 1.8] with Lemma 3.8. ∎
@@
    $\psi$ is an $(S,F,\varepsilon)$-approximately central equivariant strongly order zero map.
@@
The proof is essentially the same as [43, Lemma 4.5]; we simply replace [43, Lemma 2.9] with Lemma 3.9. ∎
When $A$ is separable, we characterize the weak tracial Rokhlin property with comparison in terms of central sequences. We begin with a definition.
@@
By (7), (8), and (10), we can choose $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, we have $\|\gamma_{n}(1)\gamma_{n}(f_{1}f_{2})-\gamma_{n}(f_{1})\gamma_{n}(f_{2})\|<\varepsilon$ for all $f_{1},f_{2}\in S$ and $\|\gamma_{n}(f)a-a\gamma_{n}(f)\|<\varepsilon$ for all $a\in F$ and $f\in S$.
Since $1-d\precsim_{\alpha-p.w.-A}x$, we can choose $n_{1}\in\mathbb{N}$ such that for all $n\geq n_{1}$, we have $(1-h_{n}-\varepsilon)_{+}\precsim_{A}x$. Similarly, since $1-d\precsim_{p.w.-A^{\alpha}}y$, we can choose $n_{2}\in\mathbb{N}$ such that for all $n\geq n_{2}$, we have $(1-h_{n}-\varepsilon)_{+}\precsim_{A^{\alpha}}y$. Similarly, since $1-d\precsim_{p.w.-A^{\alpha}}d$, we can choose $n_{3}\in\mathbb{N}$ such that for all $n\geq n_{3}$, we have $(1-h_{n}-\varepsilon)_{+}\precsim_{A^{\alpha}}h_{n}$. Also, we can choose $n_{4}\in\mathbb{N}$ such that for all $n\geq n_{4}$, we have $\|h_{n}xh_{n}\|>1-\varepsilon$. Set $N=\max(n_{0},n_{1},n_{2},n_{3},n_{4})$, choose $\psi=\gamma_{N}$ and $d=h_{N}$, then all conditions are satisfied.
@@
    $\psi_{n}$ is an ($S_{n}$$F_{n}$$\frac{1}{n}$)-approximately central strongly order zero map.
@@
We now introduce an “averaging process” introduced by Gardella in [17], which allows us to transfer certain properties from the original algebra to the fixed point algebra.
@@
[17, Lemma 4.2] Let $A$ be a unital C*-algebra, let $G$ be a compact group, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be a continuous action. Let $\varphi\colon C(G,A)\to C(G,A)$ be the averaging process defined above, and let $\gamma\colon G\to\mathrm{Aut}(C(G,A))$ be the diagonal action. Given $\varepsilon>0$ and a compact set $F\subseteq C(G,A)$, there exist $\delta>0$, a finite set $K\subseteq G$, and continuous functions $f_{k}\in C(G)$ for $k\in K$ such that
@@
The following proof references [17, Proposition 4.3].